OpenStudy (anonymous):

Solve using the square root method. x2 = 64 {8} {-8, 8} {-16, 16} {16}

OpenStudy (anonymous):

This should be asked in the Math section :] Square root method: Take the square root of both sides \[\sqrt{x^2} = \sqrt{64}\]\[x = \pm8\] 64 has 2 square roots: -8 and 8 because if you multiply (-8) by (-8) you get 64 and if you multiply (8) by (8) you get 64.

OpenStudy (anonymous):

Remember that a negative number times a negative number gives a positive number. Ex. (-2) * (-3) = 6

OpenStudy (anonymous):

It's 8. Haha. Cause 8x8= 64. And so the square root of 64 would be 8! Because 8^2=64 ! :D Heh. Hope I helped. c: